if 11 added to a number the sum is three times another number but if 9 is subtracted from the number the difference is half the other what are the two numbers ?

Answer by mananth(12270)

numbers be x, y
11+x=3y
x-3y=-11..............1
x-9 = 1/2 y
2(x-9)=y
2x-18=y
y-2x=-18..............2
...
x-3y=-11 .............1
-2x+y =-18 .............2

multiply (1)by 2
Multiply (2) by 1
2x-6y =-22
-2x+y =-18

Add the two equations

-5y=-40

/-5

y=8
Plug value of y in
x-3y=-11
x-24=-11
x=13